OpenStudy (anonymous):

Simplify the problem below. -5 × 9 -21 + 6

OpenStudy (anonymous):

2 equations?

OpenStudy (anonymous):

no. one

OpenStudy (anonymous):

they r connected?

OpenStudy (dobby1):

what is 5 * 9

OpenStudy (anonymous):

the line wouldnt show up.

OpenStudy (dobby1):

ok first what is 5*9

OpenStudy (anonymous):

Oh the equation is -5x9/-21 + 6?

OpenStudy (anonymous):

yes. and 45

OpenStudy (anonymous):

answer is 3 ask me for explanation if u desire

OpenStudy (mathstudent55):

This? \( \dfrac{-5 \times 9}{-21 + 6} \)

OpenStudy (anonymous):

The answer is 3

OpenStudy (dobby1):

\[\frac{ -5\times9 }{ -21+6 }\]

OpenStudy (anonymous):

-5x9/-21 + 6 = -45/-15 = 3

OpenStudy (anonymous):

thank you! i get it

OpenStudy (dobby1):

@greenlegodude57 you are sapposed to explain the steps not just give the answer

OpenStudy (anonymous):

I did explain.

OpenStudy (anonymous):

Look: -5 x 9 = -45 and -21 + 6 = -15. So you have -45/-15 and -45 / -15 = 3.